OpenStudy (anonymous):

Solve log7 (x – 8) – log7 3 = log7 21

OpenStudy (anonymous):

Try to show you work :) Whenever you stop i will help ..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x-8/3=21

OpenStudy (anonymous):

Nope. Show your work step by step to see which part is hard ?

OpenStudy (anonymous):

no this is where im stuck at

OpenStudy (anonymous):

hmm,Actually thats wrong and that mean that the previous step which lead to this was wrong..

OpenStudy (anonymous):

Btw : here is hint \[Logx=Logy\] x=y

OpenStudy (anonymous):

well all i did was put log7(x-8)over log7(3)=log7 21 then just took off the log7

OpenStudy (anonymous):

Its not \[ \Large \frac{Log_{7}(x-8)}{Log_{7}(3)}\] It is : \[ \Large Log_{7} \frac{(x-8)}{(3)}\]

OpenStudy (anonymous):

So now u have \[ \Large \frac{x-8}{3}=21\] 63=x-8 x=71

OpenStudy (anonymous):

ok but were did you get 63

OpenStudy (anonymous):

ooh neverminod i got it thanks

OpenStudy (anonymous):

21*3

OpenStudy (anonymous):

lets look over some basic properties such as log(a/b) = log a - log b and log(a*b)= log a + log b and if logx= log y then x=y just because everything is in the base of 7 we could use these properties as well so therefore from your question log7 (x – 8) – log7 3 = log7 21 let us use the first property on the left side and get log7((x-8)/3) = log 7 21and now let us use the third property to obtain (x-8)/3 = 21 and that will get us x-8 = 63 and now solve for x to obtain x = 71 :)

OpenStudy (anonymous):

thanks for the help everyone.I really needed it